Providing for Doubtful Accounts

At the end of the current year, the accounts receivable account has a debit balance of $1,835,000 and sales for the year total $25,690,000.

  1. The allowance account before adjustment has a debit balance of $12,500. Bad debt expense is estimated at 1/2 of 1% of sales.
  2. The allowance account before adjustment has a debit balance of $12,500. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $162,000.
  3. The allowance account before adjustment has a credit balance of $26,810. Bad debt expense is estimated at 3/4 of 1% of sales.
  4. The allowance account before adjustment has a credit balance of $26,810. An aging of the accounts in the customer ledger indicates estimated doubtful accounts of $171,200.

Determine the amount of the adjusting entry to provide for doubtful accounts under each of the assumptions (a through d) listed above.
Answer #1

Calculate amount of adjusting enty

a) Bad debt expense = 25690000*1%*1/2 = 128450

b) Bad debt expense = 162000+1500 = 174500

c) Bad debt expense = 25690000*1%*3/4 = 192675

d) Bad debt expense = 171200-26810 = 144390